Part 1. Suppose that you will need to purchase 25,000 bushels of Soybeans in March. The...

Part 1.

Suppose that you will need to purchase 25,000 bushels of Soybeans in March. The current spot price for Soybeans is 909 cents per bushel and the futures price for the March contract is 922 cents per bushel. One contract is for 5,000 bushels of soybeans. What position do you take in the March futures contract to hedge this purchase? (specify long or short and the number of contracts)

Part 2.

In the question above what price per bushel has you effectively locked in for the purchase of the Soybeans?
Answer #1

Since you have to purchase the Soybeans, you will go long on the Soybean futures.

Now, one contract is for 5,000 bushels, to purchase 25000 bushels, you will need 25000/5000 = 5 contracts.

(Note: Here the underlying security is same hence, the hedge ratio is one, hence, number of contracts derived assuming h = 1. Formula for number of contracts is;

N=h\times \frac{N_{a}}{Q_{a}})

Hence, to hedge your position, you will need 5 futures contracts of Soybeans.

Now, the Future price is of 922 cents, while the spot price is 909. Hence, the basis of the future contract is 922 - 909 = 13 cents

Basis of the futures contract is generally for transportation cost, storage cost and other factors.

The basis of the futures contract is generally very predictable and remains more or less the same.

Hence, the effective price of the Soybeans locked is 909 cents only.